A plastic container in semispheric form have 80kg. Determined the minimum radio that container must have to avoid fully sunk in the water, when we add a mass of 5.00 * 10^3 kg inside the container.
Answer #1

Solution:

Volume of a sphere is (4/3)πr3. Hence the volume of a semispheric container is half the volume of a sphere;

V = (4/3)πr3/2

V = (4/6)πr3

The mass of container and the load together,

m = 80 kg + 5.00*103 kg

m = 5080 kg

The buoyant force is given by,

fb = mf*g                   where mf = mass of the fluid displaced by the floating object

But when the object of weight Fg = m*g floats in the fluid then the buoyant force equals to the weight of the object;

Fg = Fb

Hence,

m*g = mf*g

It means a floating object displaces the fluid whose weight is equal to the weight of the object.

m = mf

we know density = mass/volume, hence for water, mf = σf*Vf

m = σf*Vf

where σf = density of water = 1000 kg/m3 and Vf* = volume of water that is being displaced by the object.

Vf = m/σf

Vf = 5080 kg/(1000 kg/m3)

Vf = 5.080 m3

Thus volume of water that is displaced by the floating object of mass m is 5.080 m3.

Now the plastic container (object) is just fully submerged, thus its volume V must the same as that of the displaced volume of water Vf

Hence

V = 5.080 m3.

(4/6)πr3 = 5.080 m3

r3 = 6*(5.080 m3)/(4π)

r = 3√[6*(5.080 m3)/(4*3.1416)]

r = 1.3436 m

Thus the minimum radius of the container should be 1.3436 m. If the radius greater than this then object will float.

Hence the answer is 1.3436 m